5. Real versus nominal GDP Consider a simple economy that produces two goods: pencils and oranges. The following table shows the prices and quantities of the goods over a three-year period Pencils Oranges Price (Dollars per orange) 2 4 4 Price Year 2012 2013 2014 Quantity (Number of pencils) 145 165 110 Quantity (Number of oranges) 195 225 165 (Dollars per pencil) Use the information from the preceding table to fill in the following table Nominal GDP Real GDP (Dollars) (Base year 2012, dollars) GDP Deflator Year 2012 2013 2014 From 2013 to 2014, nominal GDP and real GDP decreased The inflation rate in 2014 was increased5. Real versus nominal GDP Consider a simple economy that produces two goods: pencils and oranges. The following table shows the prices and quantities of the goods over a three-year period Pencils Oranges Price Price Year 2012 2013 2014 Quantity (Number of pencils) 145 165 110 Quantity (Number of oranges) 195 225 165 (Dollars per pencil) (Dollars per orange) 2 4 4 Use the information from the preceding table to fill in the following table Nominal GDP Real GDP Year(Dollars) 2012 2013 2014 (Base year 2012, dollars) GDP Deflator From 2013 to 2014, nominal GDP and real GDP decreased The inflation rate in 2014 was increasedPencils Oranges Price Quantityy Price Quantity Year (Dollars per penci) (Number of pencils) (Dollars per orange) (Number of oranges) 2012 2013 2014 145 165 110 195 225 165 2 4 4 Use the information from the preceding table to fill in the following table Nominal GDP Real GDP Year(Dollars) 2012 2013 2014 (Base year 2012, dollars) GDP Deflator From 2013 to 2014, nominal GDP ,and real GDP The inflation rate in 2014 was -12.5% Why is real GDP a more accur O Real GDP does not i O Nominal GDP is adj Real GDP is not inf of an economys production than nominal GDP? 0.1% value of intermediate goods and services, but nominal GDP does e effects of inflation or deflation, whereas real GDP is not rice changes, but nominal GDP is 12.5% 88.9% 112.590Why is real GDP a more accurate measure of an economys production than nominal GDP? O Real GDP does not include the value of intermediate goods and services, but nominal GDP does. O Nominal GDP is adjusted for the effects of inflation or deflation, whereas real GDP is not. O Real GDP is not influenced by price changes, but nominal GDP is.
Answer #1

For 2012, nominal GDP = 1*145+2*195 = $535

For 2012, real GDP = 1*145+2*195 = $535

since 2012 is the base year, then nominal GDP, will be equal to real GDP.

GDP deflator = (nominal GDP/real GDP)*100

For 2012, GDP deflator = (535/535)*100 = 100

For 2013, nominal GDP = 2*165+4*225 = $1230

For 2013, real GDP = 1*165 + 2*225 = $615

For 2013, GDP deflator = (1230/615)*100 = 200

For 2014, nominal GDP = 3*110 + 4*165 = $990

For 2014, real GDP = 1*110 + 2*165 = $440

For 2014, GDP deflator = (990/440)*100 = 225

So, following table is created

Nominal GDP Real GDP Year(Dollars) (Base year 2012, dollars) GDP Deflator 2012 535 2013 1230 2014 990 535 615 440 100 200 225

Correct Answer:

Decreased

Decreased

Inflation rate in 2014 = (225-200)/200 = 12.5%

Correct Answer:

C

Real GDP is price adjusted GDP at the base year prices. But, nominal GDP is not adjusted.